in 1984, my father was 3 times as old as me. in 2000, my father was twice as old as me. what year was my father born, and what year was I born.
Answer by stanbon(75887)

in 1984, my father was 3 times as old as me. in 2000, my father was twice as old as me. what year was my father born, and what year was I born.
-----------------
Equations:
1984 Equation:
f = 3m
-----
2000 Equation:
f+16 = 2(m+16)
------------------------
3m + 16 = 2m + 32
m = 16
f = 3m = 48 (father's age in 1984)
----
Father born in 1984-48 = 1936
"Me" born in 1984 - 16 = 1968
============
Cheers,
Stan H.
----------